An infinite number of charges each equal to Q are placed along the x axis at x=1, x=2,  x=4, x=8, …...and so on. Find potential and electric field at x=0 due to the set of charges.
 

Grade:12th pass

1 Answers

Arun
5 years ago
 

potential =kq/r in terms of magnitude

in this question the net potential is sum of all potentials

kq+kq/2+kq/4.........infinity

it is an infinite gp

net potential=a/1-r=kq/1-1/2

2kq

similarly magnitude of net field=summation of kq/r2  

kq+kq/4+kq/16......infinity

=kq/1-1/4

=4/3kq